# A note on generation and descent for derived categories of noncommutative schemes

### v3

@@ Abstract.
This work demonstrates classical generation is preserved by the derived pushforward along the canonical morphism of a noncommutative scheme to its underlying scheme. There are intriguing examples illustrating this phenomenon, particularly from noncommutative resolutions, categorical resolutions, and homological projective duality. Additionally, we establish that the Krull dimension of a variety over a field is a lower bound for the Rouquier dimension of the bounded derived category associated with a noncommutative scheme on it. This is an extension of a classical result of Rouquier to the noncommutative context.
@@
Derived categories, strong generation, Rouquier dimension, noncommutative scheme, noncommutative crepant resolutions, categorical resolution of singularities, homological projective duality, Azumaya algebra
@@ 1. Introduction
In this note, we examine the behavior of generation along the derived pushforward of the canonical morphism from a noncommutative scheme to its underlying scheme. As a result, our investigation introduces new methods coming from noncommutative homological techniques for establishing upper bounds on the Rouquier dimension within triangulated categories relevant to the commutative world.
@@
There has been recent advancements that illuminate sufficient conditions under which $D^{b}_{\operatorname{coh}}(X)$ possesses such objects: for strong generators, any Noetherian quasi-excellent separated scheme of finite Krull dimension [Aok21]; for classical generators, any Noetherian $J\textrm{-}2$ scheme [ELS20]. Albeit being aware of the existence of such objects, explicit descriptions are limited to a few instances: smooth quasi-projective schemes over a field [Rou08], Frobenius pushforwards on a compact generator for Noetherian schemes of prime characteristic [BIL+23], singular varieties that admit a resolution of singularities [Lan23, DL24], and objects arising from module categories [DLT23].
@@
Let’s now focus on understanding the Rouquier dimension of $D^{b}_{\operatorname{coh}}(X)$ in the context of a singular variety $X$. In [Lan23, DL24], it was demonstrated that if $\pi\colon\widetilde{X}\to X$ is a resolution of singularities and $G$ is a strong generator for $D^{b}_{\operatorname{coh}}(\widetilde{X})$, then $\mathbb{R}\pi_{\ast}G$ is a strong generator for $D^{b}_{\operatorname{coh}}(X)$. This provides effective methods for explicitly describing strong generators although the associated generation time remains difficult to establish.
@@
Recent attention has been drawn to the case where $X$ is a projective curve [BD11, BDG17b, HH23, BD17] with significant progress made in establishing upper bounds on the Rouquier dimension of its bounded derived category. In particular, [BDG17b] identified an essentially surjective functor $D^{b}_{\operatorname{coh}}(\mathbb{X})\to D^{b}_{\operatorname{coh}}(X)$, where $\mathbb{X}$ is a noncommutative projective curve, yielding a categorical resolution of $X$. In the context of this work, a noncommutative scheme is defined as a pair $(X,\mathcal{A})$, where $X$ is a Noetherian scheme, and $\mathcal{A}$ is a coherent $\mathcal{A}$-algebra [YZ06, BDG17a]. Additional background information can be found in Section 3.1. For alternative approaches to noncommutative algebraic geometry, refer to [Orl16, KR00, Lau03, BRS+16].
@@
In a closely related context, [ELS20] demonstrated that the bounded derived category of coherent $\mathcal{A}$-modules, denoted $D^{b}_{\operatorname{coh}}(\mathcal{A})$, admits a classical generator whenever $\mathcal{A}$ is a coherent $\mathcal{O}_{X}$-algebra over a Noetherian $J\textrm{-}2$ scheme. Our work is motivated by two key aspects: the preservation of generation along a derived pushforward of a proper surjective morphism and the geometry of a noncommutative scheme. This represents a blend of ideas from [Lan23, DL24, ELS20, BDG17a] and leads us to our first result: an investigation into the preservation of classical generators through the derived pushforward along the canonical morphism from a noncommutative scheme to its underlying scheme. The following is Theorem 3.1.
@@
Let $X$ be a Noetherian $J\textrm{-}2$ scheme of finite Krull dimension. Suppose $\mathcal{A}$ is a coherent $\mathcal{O}_{X}$-algebra with full support and structure map $\pi\colon\mathcal{O}_{X}\to\mathcal{A}$. If $G$ is a classical generator for $D^{b}_{\operatorname{coh}}(\mathcal{A})$, then $\mathbb{R}\pi_{\ast}G$ is a classical generator for $D^{b}_{\operatorname{coh}}(X)$.
There are interesting examples where Theorem A can be studied, as explored in more detail in Section 4. In particular, these include noncommutative resolutions (NCR), categorical resolution of singularities, and homological projective duality. These examples not only showcase interesting applications of noncommutative schemes but also provide valuable strategies offering upper bounds for Rouquier dimension within the commutative setting. Furthermore, we introduce a construction that yields examples of NCRs, where the obtained bounds indicate that the Rouquier dimension is no larger than the global dimension of a specific endomorphism ring. This is illustrated in Lemma 4.2.3 and discussed further in Remark 4.2.4. The subsequent Corollary 3.2 encapsulates the results presented below.
@@
In Corollary B, we may relax the requirement for the classical generator $G$ to be a bounded complex of $(S,S)$-modules, as mentioned in Remark 3.2.10. For those with a geometric inclination, Theorem 3.1 yields an intriguing consequence that is inspired by the exploration of derived categories of Azumaya algebras in [Kuz06], cf. Corollary 3.3.
Corollary C.
Let $X$ be an embeddable variety over a field. Suppose $\mathcal{A}$ is an Azumaya algebra on $X$. If $G$ is a classical generator for $D^{b}_{\operatorname{coh}}(\mathcal{A})$, then
$\dim D^{b}_{\operatorname{coh}}(X)\leq\operatorname{level}^{\pi_{\ast}G}(\mathcal{O}_{X})\cdot(\dim D^{b}_{\operatorname{coh}}(\mathcal{A})+1)-1.$
Initially, we had a similar result for Azumaya algebras over an embeddable variety utilizing homotopy colimits and a form of generation in the unbounded derived category, but the bound was far less sharp. This is discussed in Remark 3.2.12. Our last result provides lower bounds for the Rouquier dimension of the bounded derived category of a coherent $\mathcal{O}_{X}$-algebras on an integral scheme of finite type over a field, cf. Proposition 3.4. It is interesting to note that a similar bound holds for varieties over a field, see [Rou08, Proposition 7.16]. The following result is a special case of a more general statement we prove later, cf. Proposition 3.2.14.
Proposition D.
Let $X$ be an integral scheme of finite type over a field. If $\mathcal{A}$ is a coherent $\mathcal{O}_{X}$-algebra with full support, then the Rouquier dimension of $\dim D^{b}_{\operatorname{coh}}(\mathcal{A})$ is at least $\dim X$.
@@ Acknowledgements.
The authors would like to thank Ryo Takahashi for clarifying Lemma 3.2.8 and Timothy De Deyn for pointing out typos in an earlier draft. Souvik Dey was partially supported by the Charles University Research Center program No. UNCE/24/SCI/022 and a grant GA CR 23-05148S from the Czech Science Foundation. Moreover, the third author would like to personally thank Noah Olander for discussions on homotopy colimits and its relation to big generation, cf. Remark 2.10 for details. In our earlier draft, it was these techniques that were utilized to obtain Corollary 3.3, but recent work of [LO24] have allowed us to obtain sharper bounds for the same statement with a different strategy in the proof.
@@ 2. Generation
This section revisits two crucial notions of generation in triangulated categories, with a primary focus on those categories constructed from quasi-coherent sheaves on a Noetherian scheme. For foundational information on triangulated categories and generation, please consult [Kra22, Huy06, BvdB03, Rou08, Nee21, ABIM10, OŠ12a, Let21]. We fix a triangulated category $\mathcal{T}$ with shift functor $[1]\colon\mathcal{T}\to\mathcal{T}$.
Convention 2.1.
Unless otherwise stated, Noetherian ring will mean a right Noetherian and a module will mean a right module.
Notation 2.2.
@@
By abuse of notation, we let $D^{b}_{\operatorname{coh}}(R)$ denote $D^{b}_{\operatorname{coh}}(\operatorname{Spec}(R))$ if $R$ is a commutative Noetherian ring.
@@
    $\langle\mathcal{S}\rangle_{0}$ is the full subcategory consisting of all objects that are isomorphic to the zero object
@@
    $\langle\mathcal{S}\rangle_{1}$ is the full subcategory containing $G$ closed under shifts and retracts of finite coproducts
@@
From this construction, it may be verified that there is an exhaustive filtration:
@@
Let us assume that our triangulated category $\mathcal{T}$ admits all small coproducts. It should be noted that the following discussion is utilized only in Corollary 3.3. Fix an object $G$ of $\mathcal{T}$. Let $\overline{\langle G\rangle}$ denote the smallest triangulated subcategory in $\mathcal{T}$ containing $G$ that is closed under small coproducts and retracts. There exists an ascending chain of additive subcategories in $\overline{\langle G\rangle}$:
$\overline{\langle G\rangle}_{0}\subseteq\overline{\langle G\rangle}_{1}\subseteq\cdots\subseteq\overline{\langle G\rangle}$
where each $\overline{\langle G\rangle}_{i}$ is defined as follows
  1. (1)
    $\overline{\langle G\rangle}_{0}$ is the full subcategory consisting of all objects isomorphic to the zero object
  2. (2)
    $\overline{\langle G\rangle}_{1}$ is the smallest full subcategory containing $G$ closed under shifts and retracts of small coproducts
  3. (3)
    For any integer $n\geq 2$, $\overline{\langle G\rangle}_{n}$ is the full subcategory of objects which are retracts of an object $E$ appearing in a distinguished triangle
    $A\to E\to B\to A[1]$
    where $A\in\overline{\langle G\rangle}_{n-1}$ and $B\in\overline{\langle G\rangle}_{1}$.
@@
Let $X$ be a Noetherian scheme. For any $G\in D^{b}_{\operatorname{coh}}(X)$ and integer $n\geq 0$, the following can be deduced from the proofs of Lemma 2.4-2.6 in [Nee21]:
$\overline{\langle G\rangle}_{n}\cap D^{b}_{\operatorname{coh}}(X)=\langle G\rangle_{n}.$
Consequently, if $\overline{\langle G\rangle}_{n}=D_{\operatorname{Qcoh}}(X)$, then $D^{b}_{\operatorname{coh}}(X)=\langle G\rangle_{n}$.
Remark 2.10.
(Olander) Let $X$ be a Noetherian scheme, and choose $E\in D^{b}_{\operatorname{coh}}(X)$. There exists a system $(P_{n},f_{n})$ in $D_{\operatorname{Qcoh}}(X)$, where $P_{n}\in\operatorname{perf}X$ and a distinguished triangle
[…]
where $f$ is given by $1-f_{n}$ in degree $n$, cf. [Sta23, 0DJN]. Hence from Remark 2.9 it follows that if $\operatorname{Perf}(X)\subseteq\langle G\rangle_{n}$, then $D^{b}_{\operatorname{coh}}(X)=\langle G\rangle_{2n}$.
3. Derived descent
In this section, we investigate the preservation of classical generators through the derived pushforward along the canonical morphism from a noncommutative scheme to its underlying scheme. The subsequent is our main result.
Theorem 3.1.
Let $X$ be a Noetherian $J\textrm{-}2$ scheme of finite Krull dimension. Suppose $\mathcal{A}$ is a coherent $\mathcal{O}_{X}$-algebra with full support and structure map $\pi\colon\mathcal{O}_{X}\to\mathcal{A}$. If $G$ is classical generator for $D^{b}_{\operatorname{coh}}(\mathcal{A})$, then $\mathbb{R}\pi_{\ast}G$ is a classical generator for $D^{b}_{\operatorname{coh}}(X)$.
The application of Theorem 3.1 yields immediate consequences, each with distinct algebraic and geometric flavors. These consequences offer strategies coming from noncommutative objects, providing upper bounds for Rouquier dimension within the commutative setting.
Corollary 3.2.
In Corollary 3.2, our requirement for the classical generator $G$ to be a bounded complex of $(S,S)$-modules is relatively mild and can be further relaxed (see Remark 3.2.10). Next up, we present a purely geometric application inspired by the functorial properties explored for Azumaya varieties and their bounded derived categories in [Kuz06].
Corollary 3.3.
Let $X$ be an embeddable variety over a field. Suppose $\mathcal{A}$ is an Azumaya algebra on $X$. If $G$ is a classical generator for $D^{b}_{\operatorname{coh}}(\mathcal{A})$, then
$\dim D^{b}_{\operatorname{coh}}(X)\leq\operatorname{level}^{\pi_{\ast}G}(\mathcal{O}_{X})\cdot(\dim D^{b}_{\operatorname{coh}}(\mathcal{A})+1)-1.$
For a less sharp bound compared to Corollary 3.3, but whose techniques are interesting in their own right, please refer to Remark 3.2.12 where we utilize homotopy colimits and a form of generation in the unbounded derived category. The last result provides lower bounds for the Rouquier dimension of the bounded derived category of a coherent $\mathcal{O}_{X}$-algebras on an integral scheme of finite type over a field.
Proposition 3.4.
Let $X$ be an integral scheme of finite type over a field. If $\mathcal{A}$ is a coherent $\mathcal{O}_{X}$-algebra with full support, then the Rouquier dimension of $\dim D^{b}_{\operatorname{coh}}(\mathcal{A})$ is at least $\dim X$.
Before delving into the proofs of our results, we take a brief detour to explore the essential background and machinery related to noncommutative schemes.
3.1. Noncommutative schemes
This section draws directly on content found in [YZ06, ELS20, BDG17a]. For further details, the reader is encouraged to refer to these sources; especially [ELS20, Section 4] where background is sourced from. Consider a Noetherian scheme $X$. Let $\mathcal{A}$ be an $\mathcal{O}_{X}$-algebra. The notions of coherent and quasi-coherent $\mathcal{A}$-modules are defined in the obvious way. We denote the respective full subcategories of quasi-coherent and coherent sheaves in the category $\operatorname{Mod}\mathcal{A}$ of $\mathcal{A}$-modules by $\operatorname{Qcoh}\mathcal{A}$ and $\operatorname{coh}\mathcal{A}$. Note that $\operatorname{coh}\mathcal{A}$ is a full abelian subcategory of $\operatorname{Mod}\mathcal{A}$, which is also abelian. Since $\mathcal{A}$ is an $\mathcal{O}_{X}$-algebra, there exists a corresponding structure map $\pi\colon\mathcal{O}_{X}\to\mathcal{A}$. If $E\in\operatorname{Mod}\mathcal{A}$, we can view $E$ as a sheaf of $\mathcal{O}_{X}$-modules by locally restricting scalars via $\pi$. We denote the resulting sheaf of $\mathcal{O}_{X}$-modules as $\pi_{\ast}E$.
Remark 3.1.1.
@@ 3.1. Noncommutative schemes
This pair is referred to as the canonical morphism. To simplify notation, we abuse notation to denote it as $\pi$.
Example 3.1.4.
Let $X$ be a variety over a field. Consider a sheaf of Azumaya algebras $\mathcal{A}$ on $X$, i.e., a locally finite free sheaf on $X$ that is étale locally a matrix algebra. This sheaf corresponds to an element in the Brauer group of $X$. For a treatment of the functorial properties of the noncommutative scheme $(X,\mathcal{A})$ and its bounded derived category, refer to [Kuz06, Appendix D]. The pair $(X,\mathcal{A})$ is called a Azumaya variety.
Remark 3.1.5.
Let $\pi\colon Y\to X$ be a morphism of schemes, and set $\mathcal{B}:=\pi^{\ast}\mathcal{A}$
@@
    Remark 3.1.1 tells us $\mathcal{B}:=\pi^{\ast}\mathcal{A}$ is an $\mathcal{O}_{Y}$-algebra. Moreover, if $\mathcal{A}$ is (quasi-)coherent, then so is $\mathcal{B}$.
@@
    If $\pi^{\ast}$ and $\pi_{\ast}$ are defined in the usual manner, then we have an adjunction which commutes with the forgetful functor to modules over the underlying schemes:
[…]
@@
    If $\mathcal{A}$ is coherent, then $\pi^{\ast}$ restricts to $\pi^{\ast}\colon\operatorname{coh}\mathcal{A}\to\operatorname{coh}\mathcal{B}$.
@@
where $Z:=X\setminus Y$. This is [ELS20, Theorem 4.4].
Example 3.1.7.
@@
Remark 3.1.8.
@@
3.2. Results
To exhibit Theorem 3.1, we start by checking for exactness of the canonical morphism of a noncommutative scheme to its underlying scheme. The following statement might be clear to experts, but we include it for the sake of coherence.
Lemma 3.2.1.
Let $X$ be a Noetherian scheme. If $\mathcal{A}$ is a coherent $\mathcal{O}_{X}$-algebra with structure map $\pi\colon\mathcal{O}_{X}\to\mathcal{A}$, then there exists an exact functor $\pi_{\ast}\colon\operatorname{coh}\mathcal{A}\to\operatorname{coh}X$.

### v4

# Preservation for generation along the structure morphism of coherent algebras over a scheme

@@ Abstract.
This work demonstrates classical generation is preserved by the derived pushforward along the structure morphism of a noncommutative coherent algebra to its underlying scheme. Additionally, we establish that the Krull dimension of a variety over a field is a lower bound for the Rouquier dimension of the bounded derived category associated with a noncommutative coherent algebra on it. This is an extension of a classical result of Rouquier to the noncommutative context.
@@
Derived categories, Rouquier dimension, strong generators, noncommutative algebraic geometry, coherent algebras
@@ 1. Introduction
In this note, we examine the behavior of generation along the derived pushforward of the structure morphism from a noncommutative coherent algebra to its underlying scheme. As a result, our investigation introduces new methods coming from noncommutative homological techniques for establishing upper bounds on the Rouquier dimension within triangulated categories relevant to the commutative world.
@@
There has been recent advancements that illuminate sufficient conditions under which $D^{b}_{\operatorname{coh}}(X)$ possesses such objects: for strong generators, any Noetherian quasi-excellent separated scheme of finite Krull dimension [Aok21]; for classical generators, any Noetherian $J\textrm{-}2$ scheme [ELS20] or Notherian schemes whose closed subschemes admit open regular locus locus [DL24a]. Albeit being aware of the existence of such objects, explicit descriptions are limited to a few instances: smooth quasi-projective schemes over a field [Rou08], Frobenius pushforwards on a compact generator for Noetherian schemes of prime characteristic [BIL+23], singular varieties that admit a resolution of singularities [Lan24, DL24b], and objects arising from module categories [DLT23].
@@
Let’s now focus on understanding the Rouquier dimension of $D^{b}_{\operatorname{coh}}(X)$ in the context of a singular variety $X$. In [Lan24, DL24b], it was demonstrated that if $\pi\colon\widetilde{X}\to X$ is a resolution of singularities and $G$ is a strong generator for $D^{b}_{\operatorname{coh}}(\widetilde{X})$, then $\mathbb{R}\pi_{\ast}G$ is a strong generator for $D^{b}_{\operatorname{coh}}(X)$. This provides effective methods for explicitly describing strong generators although the associated generation time remains difficult to establish.
@@
Recent attention has been drawn to the case where $X$ is a projective curve [BD11, BDG17b, HH23, BDG17a] with significant progress made in establishing upper bounds on the Rouquier dimension of its bounded derived category. In particular, [BDG17b] identified an essentially surjective functor $D^{b}_{\operatorname{coh}}(\mathbb{X})\to D^{b}_{\operatorname{coh}}(X)$, where $\mathbb{X}$ is a noncommutative projective curve, yielding a categorical resolution of $X$. In the context of this work, a noncommutative scheme is defined as a pair $(X,\mathcal{A})$, where $X$ is a Noetherian scheme, and $\mathcal{A}$ is a coherent $\mathcal{A}$-algebra [YZ06, BDG17a, DLM24b]. Additional background information can be found in Section 3. There alternative approaches to ‘noncommutative algebraic geometry’ for which we refer the reader to [Orl16, KR00, Lau03, BRS+16].
@@
In a closely related context, [ELS20] demonstrated that the bounded derived category of coherent $\mathcal{A}$-modules, denoted $D^{b}_{\operatorname{coh}}(\mathcal{A})$, admits a classical generator whenever $\mathcal{A}$ is a coherent $\mathcal{O}_{X}$-algebra over a Noetherian $J\textrm{-}2$ scheme (see Remark 2.6 for details on terminology). There has been recent attention towards the existence of strong generators in this noncommutative setting [DLM24b].
Our work is motivated by two key aspects: the preservation of generation along a derived pushforward of a proper surjective morphism and the geometry of a noncommutative scheme. This represents a blend of ideas from [Lan24, DL24b, ELS20, BDG17a] and leads us to our first result: an investigation into the preservation of classical generators through the derived pushforward along the structure morphism from a noncommutative scheme to its underlying scheme. This brings attention to our first result.
@@
(see Theorem 4.4) Let $X$ be a Noetherian $J\textrm{-}2$ scheme of finite Krull dimension. Suppose $\mathcal{A}$ is a coherent $\mathcal{O}_{X}$-algebra with full support and canonical map $\pi\colon\mathcal{O}_{X}\to\mathcal{A}$. If $G$ is classical generator for $D^{b}_{\operatorname{coh}}(\mathcal{A})$, then $\mathbb{R}\pi_{\ast}G$ is a classical generator for $D^{b}_{\operatorname{coh}}(X)$.
There are interesting examples where Theorem A can be applied. This includes categorical resolutions [^1] See [KL15, Lun10, Kuz07]., homological projective duality [^2] See [Kuz06] and more recently [Per19, KP21]., and noncommutative crepant resolutions [^3] See [VdB04, SVdB08, IW13, DFI15].. These are special cases of noncommutative schemes which could be leveraged for studying generation in the commutative setting.
The next result gives a way to make this precise in mild situations.
@@
(see Corollary 4.7) Let $R$ be a commutative Noetherian $J\textrm{-}2$ ring of finite Krull dimension and $\pi\colon R\to S$ a finite ring morphism such that $\pi_{\ast}S$ has full support as an $R$-module. If $G$ is a classical generator for $D^{b}_{\operatorname{coh}}(S)$ which is a bounded complex of $(S,S)$-bimodules, then
@@
Another interesting problem is to provide lower bounds for the Rouquier dimension of a triangulated category. Our last result provides lower bounds for the Rouquier dimension of the bounded derived category of a coherent $\mathcal{O}_{X}$-algebras on a general class of schemes.
Theorem C.
(see Theorem 4.8) Let $X$ be a scheme of finite Krull dimension which is integral, Jacobson, catenary, $J\textrm{-}2$, and Noetherian. If $\mathcal{A}$ is a coherent $\mathcal{O}_{X}$-algebra with full support, then the Rouquier dimension of $D^{b}_{\operatorname{coh}}(\mathcal{A})$ is at least the Krull dimension of $X$.
Theorem 4.8 is applicable to many cases of interest. This includes not only coherent algebras with full support over a variety, but also for such algebras over rings of mixed characteristic (i.e over $\mathbb{Z}$). It is worthwhile to note that in the commutative setting a similar bound holds for varieties over a field, see [Rou08, Proposition 7.16].
@@ Acknowledgements.
The authors would like to thank Ryo Takahashi for clarifying Lemma 4.6, Timothy De Deyn for comments in an earlier draft, and the anonymous referee for many helpful suggestions. Souvik Dey was partially supported by the Charles University Research Center program No. UNCE/24/SCI/022 and a grant GA CR 23-05148S from the Czech Science Foundation.
@@ 2. Generation
This section revisits two crucial notions of generation in triangulated categories, with a primary focus on those categories constructed from quasi-coherent sheaves on a Noetherian scheme. For foundational information on generation, please see [BvdB03, Rou08, ABIM10]. We fix a triangulated category $\mathcal{T}$ with shift functor $[1]\colon\mathcal{T}\to\mathcal{T}$. Unless otherwise stated, all rings considered are Noetherian, and a ‘module’ means a right module.
Notation 2.1.
@@
By abuse of notation, we let $D^{b}_{\operatorname{coh}}(R)$ denote $D^{b}_{\operatorname{coh}}(\operatorname{Spec}(R))$ if $R$ is a commutative Noetherian ring, and similarly for other associated categories.
@@
    $\langle\mathcal{S}\rangle_{0}=0$
@@
    $\langle\mathcal{S}\rangle_{1}$ is the full subcategory containing $\mathcal{S}$ closed under shifts and retracts of finite coproducts
@@
If $\mathcal{S}$ consist of a single object $G$, then we set $\langle\mathcal{S}\rangle_{n}=\langle G\rangle_{n}$.
@@
It can be seen that $\bigcup_{n=0}^{\infty}\langle\mathcal{S}\rangle_{n}$ is contained in $\langle\mathcal{S}\rangle$ because of how each $\langle\mathcal{S}\rangle_{n}$ is defined. An induction argument will show that $\bigcup_{n=0}^{\infty}\langle\mathcal{S}\rangle_{n}$ is a thick subcategory of $\mathcal{T}$ containing $\mathcal{S}$, and so, $\langle\mathcal{S}\rangle$ is a subcategory of $\bigcup_{n=0}^{\infty}\langle\mathcal{S}\rangle_{n}$. Tying this together gives us an exhaustive filtration:
@@
Remark 2.6.
Let $X$ be a Noetherian scheme. There is an interesting connection with openness of the regular locus [^4] The collection of points $p$ in $X$ such that $\mathcal{O}_{X,p}$ is regular. of $X$ and existence of classical generators for $D^{b}_{\operatorname{coh}}(X)$. Specifically, it has been shown that every closed integral subscheme $Z$ of $X$ has an open regular locus if, and only if, $D^{b}_{\operatorname{coh}}(Z)$ admits a classical generator (see [DL24a, Theorem 1.1]). This particular result was initially studied in the affine setting [IT19]. The openness of a regular locus for a scheme has be of interest outside of generation problems, and leads to the following well-studied notions (see [Sta23, Tag 07P6] and [Sta23, Tag 07R2] for details):
  1. (1)
    $X$ is J-0 if the regular locus of $\mathcal{X}$ contains a nonempty open subset of $X$
  2. (2)
    $X$ is J-1 if the regular locus of $X$ is open
  3. (3)
    $X$ is J-2 if the regular locus of $Y$ is open for every morphism $Y\to X$ that is locally of finite type.
@@
3. Noncommutative schemes
This section draws directly on content found in [YZ06, ELS20, BDG17a]. For further details, the reader is encouraged to refer to these sources; especially [ELS20, Section 4] where background is sourced from. Consider a Noetherian scheme $X$. Let $\mathcal{A}$ be an $\mathcal{O}_{X}$-algebra. The notions of coherent and quasi-coherent $\mathcal{A}$-modules are defined in the obvious way. We denote the respective full subcategories of quasi-coherent and coherent sheaves in the category $\operatorname{Mod}\mathcal{A}$ of $\mathcal{A}$-modules by $\operatorname{Qcoh}\mathcal{A}$ and $\operatorname{coh}\mathcal{A}$. Note that $\operatorname{coh}\mathcal{A}$ is a full abelian subcategory of $\operatorname{Mod}\mathcal{A}$, which is also abelian. Since $\mathcal{A}$ is an $\mathcal{O}_{X}$-algebra, there exists a corresponding canonical map $\pi\colon\mathcal{O}_{X}\to\mathcal{A}$. If $E\in\operatorname{Mod}\mathcal{A}$, we can view $E$ as a sheaf of $\mathcal{O}_{X}$-modules by locally restricting scalars via $\pi$. We denote the resulting sheaf of $\mathcal{O}_{X}$-modules as $\pi_{\ast}E$.
@@ 3. Noncommutative schemes
    This pair is referred to as the structure morphism, and we abuse notation to denote it as $\pi$.
@@
    Let $i\colon Z\to X$ be a closed immersion $i\colon Z\to X$. This induces a morphism of noncommutative schemes $(Z,i^{\ast}\mathcal{A})\to(X,\mathcal{A})$, which we abuse notation and denote by $i$ as well. There exists a commutative diagram of noncommutative schemes:
    ${{(Z,i^{\ast}\mathcal{A})}}$${{(X,\mathcal{A})}}$${Z}$${X.}$$\scriptstyle{i}$$\scriptstyle{\theta}$$\scriptstyle{\pi}$$\scriptstyle{i}$
    Here, $\theta$ is the structure morphism $\mathcal{O}_{Z}\to j^{\ast}\mathcal{A}$.
Remark 3.4.
Let $f\colon Y\to X$ be a morphism of schemes, and set $\mathcal{B}:=f^{\ast}\mathcal{A}$
@@
    Remark 3.1 tells us $\mathcal{B}$ is an $\mathcal{O}_{Y}$-algebra. Moreover, if $\mathcal{A}$ is (quasi-)coherent, then so is $\mathcal{B}$.
  2. (2)
    There are the usual adjunctions $f^{\ast}\colon\operatorname{Mod}\mathcal{A}\to\operatorname{Mod}\mathcal{B}$ and $f_{\ast}\colon\operatorname{Mod}\mathcal{B}\to\operatorname{Mod}\mathcal{A}$ which are defined in the usual way. Note that $f^{\ast}$ is left adjoint to $f_{\ast}$. These functors commutes with the forgetful functor to modules over the underlying schemes.
@@
    If $\mathcal{A}$ is coherent, then $f^{\ast}$ restricts to $f^{\ast}\colon\operatorname{coh}\mathcal{A}\to\operatorname{coh}\mathcal{B}$.
@@
where $Z:=X\setminus Y$. Note that the exact functor $j^{\ast}\colon D^{b}_{\operatorname{coh}}(\mathcal{A})\to D^{b}_{\operatorname{coh}}(j^{\ast}\mathcal{A})$ is a small abuse of notation, but it is defined by applying $j^{\ast}$ to each component of a complex in $D^{b}_{\operatorname{coh}}(\mathcal{A})$. This is [ELS20, Theorem 4.4].
Example 3.6.
  1. (1)
@@
  2. (2)
    Let $X$ be a separated scheme of finite type over a perfect field. Then $D^{b}_{\operatorname{coh}}(\mathcal{A})$ admits a strong generator for any coherent $\mathcal{O}_{X}$-algebra, see [DLM24a, Example 3.9]. The affine setting was investigated [ELS20, Remark 2.6].
Remark 3.7.
@@
4. Descent along structure morphism
This section investigates the preservation of classical generators through the derived pushforward along the structure morphism from a noncommutative scheme to its underlying scheme. We start with a few elementary lemmas.
Lemma 4.1.
Let $X$ be a Noetherian scheme and $\mathcal{A}$ be a coherent $\mathcal{O}_{X}$-algebra. Then the canonical map $\pi\colon\mathcal{O}_{X}\to\mathcal{A}$ induces a exact functor $\pi_{\ast}\colon\operatorname{coh}\mathcal{A}\to\operatorname{coh}X$.